On May 1, 2013, Bentham Company sells office furniture for $60,000 cash. The office furniture originally cost $150,000 when purchased on January 1, 2006. Depreciation is recorded by the straight-line method over 10 years with a salvage value of $15,000. What gain should be recognized on the sale?
Answer #1

Gain on sale = Cash received - Book value

Book value = 150,000 - Accumulated Depreciation

Accumulated Depreciation =

Depreciation upto 2013 + Depreciation in 2013

= [(150,000-15,000)/10 * 7 years] + [150,000-15000)/10* 4/12]

= 94,500 + 4500

= 99,000

Gain on sale = 60,000 - (150,000-99,000)

= 60,000 - 51,000

= 9,000